11123:

Barbara draws pens randomly from a box containing 5 pens of the same shape and size. There is 1 green pen, 3 red pens, and 1 blue pen. She draws 1 red pen and then another red pen without replacing the first one. Find the probability of drawing 1 red pen followed by another red pen, and show the equation used.

11123:

@AnimeGhoul8863

AnimeGhoul8863:

The contidion lets to establish that the events are independent. Probability is the number of positive results / number of total events The probability of drawing the first red pen is 3/5. The probability of drawing the second red pen is 2/4 The combined propability is (3/5)*(2/4) = 6/20 = 3/10 = 0.30 Soooo sorry i was answering someone else

AnimeGhoul8863:

@11123

11123:

its ok

AnimeGhoul8863:

thx anyway thats ur answer hope it helps